OpenStudy (anonymous):

A ladder leans against a window in a two-story building. The window is 16 feet off the ground. The base of the ladder is 12 feet from the base of the building. How many feet long is the ladder?

OpenStudy (anonymous):

So the proper equation is 16^2 + 12^2 = L^2 So you have 256 + 144 = L^2 Thats 400 = L^2 So the ladder is 20 feet long.